### v1

# Amazon Web Service

Abstract
Large Language Models (LLMs) display striking surface fluency yet systematically fail at tasks requiring symbolic reasoning, arithmetic accuracy, and logical consistency. This paper offers a structural diagnosis of such failures, revealing a persistent gap between comprehension and competence. Through controlled experiments and architectural analysis, we demonstrate that LLMs often articulate correct principles without reliably applying them—a failure rooted not in knowledge access, but in computational execution. We term this phenomenon the computational split-brain syndrome, where instruction and action pathways are geometrically and functionally dissociated. This core limitation recurs across domains, from mathematical operations to relational inferences, and explains why model behavior remains brittle even under idealized prompting. We argue that LLMs function as powerful pattern completion engines, but lack the architectural scaffolding for principled, compositional reasoning. Our findings delineate the boundary of current LLM capabilities and motivate future models with metacognitive control, principle lifting, and structurally grounded execution. This diagnosis also clarifies why mechanistic interpretability findings may reflect training-specific pattern coordination rather than universal computational principles, and why the geometric separation between instruction and execution pathways suggests limitations in neural introspection and mechanistic analysis.
1 Introduction
When asked whether 9.9 is greater than 9.11, Claude Sonnet 4 confidently states that 9.11 is larger, explaining that “since 90 is greater than 11 in the hundredths place, 9.11 is the larger number”—while simultaneously calculating that 9.11 - 9.9 = 0.21 (incorrect). GPT-4o claims that “9.11 is greater than 9.9” in mathematical contexts, invoking software versioning to justify the confusion. Yet when asked to explain how to compare decimal numbers, both models provide flawless algorithmic descriptions: “Write the numbers one above the other, aligning the decimal points… Compare digit by digit from left to right.” Claude Sonnet 4 even works through the exact same comparison correctly as an instructional example, concluding that “9.90 $>$ 9.11.” Both models articulate decimal comparison procedures with textbook precision while failing to execute them reliably.
This disconnect reveals a fundamental limitation in Large Language Models: they exhibit comprehension without competence—a systematic dissociation where models can perfectly explain principles they cannot reliably execute. We term this phenomenon computational split-brain syndrome, drawing analogy to neurological conditions where different brain systems cannot coordinate effectively. Like patients who can verbally describe actions they cannot perform, LLMs develop geometrically separated pathways for “knowing about” procedures versus “executing” them.
This is not a limitation of scale, training data, or prompting techniques. We argue that computational split-brain syndrome stems from structural constraints in how LLMs represent symbols, learn procedures, and execute reasoning steps. This dissociation challenges fundamental assumptions about intelligence derived from human cognition, where explanatory fluency typically correlates with execution competence. LLMs systematically violate this expectation, revealing that artificial and biological intelligence may operate under fundamentally different principles. Unlike symbolic systems that bind tokens to abstract roles and apply rules over those bindings, LLMs operate as pattern completion engines optimized for next-token prediction. This architectural difference creates three interdependent constraints that prevent reliable symbolic execution despite perfect comprehension:
  • Contextual Averaging: Token embeddings encode context-weighted averages that systematically distort symbolic relationships, preventing stable domain binding. The token “9.11” retains historical event associations that interfere with mathematical computation. Such contamination is inevitable in general-purpose models but absent in narrow-domain training.
  • Architectural Impossibility: Feed-forward networks face fundamental constraints that prevent exact symbolic computation through weight configuration alone, forcing hierarchical pattern assembly rather than principled algorithmic implementation.
  • Instruction-Execution Disconnect: Next-token prediction treats algorithmic descriptions and execution traces as equivalent pattern completion tasks, creating geometrically separated representational pathways for procedural knowledge versus computational implementation.
We demonstrate that these constraints manifest consistently across two critical domains: symbolic computation (arithmetic operations) and relational reasoning (logical inference and variable binding). In both domains, models exhibit the same split-brain syndrome—fluent explanation coupled with unreliable execution—revealing this as a fundamental architectural limitation rather than a domain-specific failure.
Main Claims of This Work.
Our analysis establishes this disconnect through three interconnected claims:
  • Claim 1 (Representation): LLM token embeddings encode context-weighted averages that systematically resist automatic domain binding, preventing stable symbolic circuits across computational domains.
  • Claim 2 (Computation): Feed-forward networks face architectural impossibility of implementing exact symbolic operations through weight configuration alone, forcing them to resort to residual pattern fitting rather than implementing generalizable symbolic procedures.
  • Claim 3 (Instruction-Execution Disconnect): Next-token prediction objectives decouple algorithmic descriptions from executable behavior.
These architectural limitations manifest consistently across arithmetic and logical domains, creating the systematic dissociation between comprehension and competence we term computational split-brain syndrome.
Contributions.
Our analysis establishes computational split-brain syndrome as a unifying framework for understanding when and why LLMs fail at systematic reasoning. We provide empirical evidence for each architectural constraint through embedding analysis, layer-by-layer computation tracking, and systematic evaluation across arithmetic and logical tasks. We then examine three compensatory strategies—self-scaffolding, tool delegation, and hybrid architectures—showing how each leverages LLMs’ pattern completion strengths while working around their execution limitations, yet all require metacognitive capabilities that current architectures fundamentally lack.
Our findings suggest that mechanistic interpretability studies identifying “arithmetic circuits” or “adder neurons” may be observing sophisticated forms of pattern matching coordination that are formed path-dependently during learning, rather than genuine computational subroutines. More critically, the geometric dissociation between instruction and execution pathways raises concerns about both model self-explanations and interpretability research: models may articulate reasoning procedures through neural routes distinct from those used for actual computation, potentially misleading both self-monitoring and researcher analysis of neural activations.
These constraints appear unavoidable within the current paradigm: contextual averaging emerges inevitably from next-token prediction over diverse corpora, while predicting next tokens on internet-scale data forces FFNs to work with attention layers into hierarchical pattern assembly rather than principled computation, and instruction-execution disconnect results from treating all text as equivalent prediction targets. This suggests computational split-brain syndrome will persist across model families and scales unless addressed through fundamental architectural innovations rather than incremental improvements.
Roadmap.
Section 2 reviews foundational work on transformer computation, symbolic reasoning, and recent interpretability findings that inform our analysis. In Section 3, we investigate LLM failures in symbolic computation, focusing on arithmetic tasks that reveal unstable embeddings and residual fitting behavior. Section 4 extends this analysis to relational reasoning, demonstrating that the same architectural bottlenecks underlie LLM failures in multi-step inference, variable binding, and logical consistency. Section 5 examines compensatory approaches—self-scaffolding, tool delegation, and hybrid architectures—revealing how all three strategies converge on the same reliable metacognitive limitations that current architectures lack. Section 6 analyzes LLMs as hierarchical pattern completion engines, distinguishing between general intelligence (which LLMs achieve through sophisticated pattern matching) and generalizable intelligence (which requires systematic rule discovery and principled reasoning), and examining the performance cliffs that emerge when tasks transition from pattern recognition to genuine algorithmic discovery.
Limitations.
This analysis focuses on pretrained transformer-based LLMs trained on next-token prediction objectives over natural language corpora, without access to external tools or explicit reasoning scaffolds. Our empirical evaluation centers on specific model families (LLaMA2, Claude, GPT-4) and may not generalize to fundamentally different architectures or training paradigms. The geometric separation experiments rely on t-SNE projections which may not capture all relevant representational structure. Our theoretical analysis of FFN computational limits focuses on ReLU networks and may not apply to other activation functions. Additionally, we do not extensively evaluate recent architectural innovations like mixture-of-experts or specialized reasoning modules that might partially address the constraints we identify.
2 Related Work
This work builds on several lines of inquiry across theory, empirical analysis, architecture, and cognitive science. We organize our review around the emergence, diagnosis, and mitigation of systematic reasoning failures in LLMs, providing a unified framework for understanding when and why current architectures fall short of reliable symbolic computation.
2.1 Foundations and Theoretical Limits
Expressivity and Computational Boundaries.
Recent theoretical frameworks have clarified the fundamental boundaries of transformer expressivity. Strobl et al. (Strobl et al., 2024) provide a comprehensive survey documenting how transformers function as recognizers of formal languages, revealing that while scaffolding strategies can improve performance within bounded expressivity classes, they remain fundamentally constrained by architectural limitations rather than achieving true symbolic reasoning capabilities. The Simons Institute workshop (Simons Institute for the Theory of Computing, 2024) established that transformers act as highly parallel pattern-matching engines with limited capacity for long-range sequential computation, constraining their theoretical expressivity in structured symbolic tasks.
These formal limitations manifest practically as a persistent gap between what LLMs can articulate versus what they can reliably execute. This disconnect becomes especially apparent in arithmetic and logic tasks, where scaffolding strategies like chain-of-thought prompting (Wei et al., 2022) and reasoning-trained models like o1 and DeepSeek-R1 (OpenAI, 2024; DeepSeek-AI et al., 2025) show improved performance yet remain fundamentally compensatory rather than curative—they bypass core architectural bottlenecks rather than resolving them.
Training Dynamics and Path Dependence.
Understanding how reasoning capabilities emerge during training reveals critical insights into LLM limitations. Power et al. (Power et al., 2022) demonstrate that late-phase generalization involves fundamental reorganization of internal representations, showing how training dynamics create path-dependent artifacts that appear as reasoning capabilities but reflect statistical regularities rather than systematic algorithms. Tigges et al. (Tigges et al., 2024) extend this analysis to decoder-only LLMs of moderate scale (up to 2.8B parameters), finding that while computational circuits emerge consistently across scale, their specific implementations vary significantly during training, suggesting that apparent “reasoning circuits” may reflect training-specific pattern coordination rather than universal computational principles, though whether this holds for larger models remains an open question. Recent mechanistic analysis of factual learning provides direct evidence for these training-dependent phenomena. Zucchet et al. (2025) demonstrate that transformer knowledge acquisition follows three distinct phases: learning general statistics, a plateau phase where attention circuits develop, followed by individual-specific knowledge acquisition. Critically, they show that adding new factual knowledge rapidly corrupts existing memories stored in feed-forward layers, confirming that apparent knowledge storage reflects fragile statistical coordination rather than stable representation systems.
2.2 Empirical Failures Across Domains
Symbolic Computation Breakdowns.
Systematic evaluation reveals consistent failures in symbolic computation despite surface fluency. Srivastava et al. (Srivastava et al., 2024) quantify sharp performance drops when problem structures deviate from training distributions. Mirzadeh et al. (Mirzadeh et al., 2024) demonstrate through the GSM-Symbolic benchmark that LLMs exhibit noticeable variance when only numerical values change, with performance drops up to 65% when irrelevant clauses are added. This fragility confirms that current LLMs cannot perform genuine logical reasoning and instead replicate reasoning patterns from training data.
Dziri et al. (Dziri et al., 2023) provide comprehensive analysis showing that transformers solve compositional tasks through “linearized subgraph matching”—memorizing computation patterns rather than learning systematic algorithms. Their frequency analysis reveals that models succeed primarily when test computation subgraphs appeared frequently in training data, failing dramatically on out-of-distribution examples. Nikankin et al. (Nikankin et al., 2024) confirm this pattern-dependent behavior, showing that models rely on a “bag of heuristics” for arithmetic rather than implementing systematic algorithms. Wu et al. (Wu et al., 2025) further demonstrate that domain-specific fine-tuning can increase answer accuracy while degrading the logical coherence of chain-of-thought explanations, reinforcing the comprehension–execution split.
Multiple recent studies provide converging evidence for the execution limitation across different domains. Yang et al. (Yang et al., 2024) show that LLMs consistently outperform chain-of-thought when generating Prolog programs for external execution rather than attempting direct calculation. The Illusion of Thinking commentary (Opus and Lawsen, 2025) demonstrates similar improvements when LLMs generate Lua code for Tower of Hanoi problems rather than attempting direct move enumeration. Cheng et al. (Cheng et al., 2024) provide systematic evidence that LLMs excel at pattern recognition and code generation while struggling with deductive reasoning, with their successful framework architecturally separating symbolic code generation from symbolic code execution. This convergent pattern across arithmetic reasoning, logic puzzles, and systematic inference demonstrates that LLMs excel at extracting predicates and generating symbolic representations while requiring external systems for reliable computation—precisely the instruction-execution disconnect we formalize.
Relational and Logical Reasoning Failures.
The same architectural constraints manifest in relational reasoning. Berglund et al. (Berglund et al., 2023) identify the “reversal curse,” where models fail to infer bidirectional relations from symmetric facts due to asymmetric training exposure. Nezhurina et al. (Nezhurina et al., 2024) show similar failures in multi-step family reasoning problems that should be trivial for systematic logical processing.
A controlled biography-corpus study by Allen-Zhu et al. (Allen-Zhu and Li, 2024) reinforces the same point: a GPT-2 model trained exclusively on person–relation tuples and matching QA pairs achieves near-perfect in-distribution accuracy, yet its OOD QA accuracy on unseen entities collapses below 10% unless the pre-training data are aggressively rewritten, permuted, or translated. This highlights that relational exposure alone is insufficient for lifted reasoning; surface-form augmentation is required to abstract away from individual instances.
Li et al. (Li et al., 2024) provide systematic evaluation across inductive logic programming tasks, finding that LLMs—despite being 100,000 times larger than specialized logical reasoning models—perform dramatically worse on tasks requiring variable binding and systematic rule application. This performance gap persists even when given explicit logical structure through truth-value matrices, definitively ruling out training data scarcity as an explanation and confirming architectural limitations.
2.3 Mechanistic Explanations
Recent mechanistic interpretability research has developed sophisticated tools for tracing computational pathways within transformer models. Attribution graph studies by Lindsey et al. (Lindsey et al., 2025) demonstrate methods for mapping how models transform inputs to outputs through intermediate representational steps, while other circuit analyses reveal consistency patterns across training and scale (Tigges et al., 2024).
Representational Pathologies.
The instruction-execution disconnect operates at the representational level through systematic failures in symbolic binding. McLeish et al. (McLeish et al., 2024) demonstrate that altering input geometry alone can dramatically improve simple numerical reasoning, confirming that embedding representations constitute a significant bottleneck for symbolic manipulation. This supports our analysis that contextual averaging prevents stable domain binding required for symbolic manipulation.
Training-Order Dependencies and Pattern Storage.
Mechanistic interpretability reveals that apparent “reasoning circuits” often reflect distributed heuristics rather than principled computation. Nikankin et al. (Nikankin et al., 2024) show that arithmetic behaviour emerges from a “bag of heuristics” spread across layers instead of dedicated algorithmic modules. Tigges et al. (Tigges et al., 2024) further demonstrate that, even when a high-level algorithm seems stable, the specific attention heads instantiating it drift throughout training, underscoring strong path dependence. Ye et al. (Ye et al., 2025) train a 124M-parameter GPT-2 exclusively on a procedurally generated grade-school math curriculum and uncover a linearly decodable dependency-graph representation that tracks which intermediate quantities depend on which others. We interpret these coherent circuits as artifacts of an unnaturally homogeneous training path; they dissolve once heterogeneous text is introduced, reinforcing our claim that stable algorithmic modules are fragile in realistic open-domain settings.
Reasoning Faithfulness and Post-Hoc Construction.
Comprehensive analysis of reasoning faithfulness reveals fundamental limitations in how LLMs construct explanations. Plaat et al. (Plaat et al., 2024) survey extensive evidence that LLM reasoning may be “post-hoc” and “constructed after a conclusion has been found,” with larger models showing less faithful reasoning. Chain-of-thought continues to work “even with invalid steps in the reasoning chain,” suggesting pattern matching rather than logical execution. This evidence strongly supports our geometric separation hypothesis—models access instructional pathways distinct from computational pathways when generating explanations.
2.4 Compensatory Strategies and Architectural Remedies
Prompt-Level and Self-Scaffolding Interventions.
Beyond standard chain-of-thought (Wei et al., 2022), researchers have proposed metacognitive variants that embed self-evaluation into the reasoning process. AbstRaL (Gao et al., 2025) trains models to first generate an abstract symbolic representation of a problem via reinforcement learning before attempting execution, substantially improving robustness under distribution shift. Self-correction frameworks such as SPOC (Zhao et al., 2025) train models to interleave solution generation with explicit verification steps, enabling dynamic revision when the internal verifier detects errors. This exemplifies the self-scaffolding pattern analyzed in Section 5.
Tool Integration and Hybrid Execution.
Tool-augmented approaches explicitly address the execution gap. Toolformer (Schick et al., 2023) enables automated tool calling, while ReAct (Yao et al., 2022) combines reasoning with external action. Yang et al. (Yang et al., 2024) provide empirical validation that this separation—LLMs for predicate extraction, external systems for computation—consistently outperforms end-to-end LLM execution across arithmetic tasks.
Architectural Modifications.
To address fundamental limitations, researchers have explored hybrid architectures incorporating explicit symbolic modules. OccamLLM (Dugan et al., 2024) and IGC (Dietz and Klakow, 2024) embed arithmetic reasoning units, while Logic-LM (Pan et al., 2023), NLM (Dong et al., 2019), and Differentiable Logic Machines (Zimmer et al., 2023) implement logical operations through structured tensors. These systems represent promising directions for true architectural solutions rather than compensatory workarounds.
2.5 Positioning and Contributions
Our work advances this literature in three crucial ways. First, while previous studies document what failures occur across domains, we provide a principled mechanistic explanation for why they occur through three interconnected architectural constraints: contextual averaging, computational impossibility, and instruction-execution disconnect. Second, rather than treating arithmetic and relational reasoning as separate problems, we demonstrate that identical underlying limitations create the computational split-brain syndrome across all symbolic domains. Third, our geometric analysis reveals that this disconnect operates at the representational level, with direct implications for interpretability research showing why mechanistic studies often surface training-dependent artifacts rather than universal computational principles.
This unified framework not only explains current limitations but predicts why certain compensatory strategies succeed while others fail, providing a foundation for developing more reliable reasoning systems that leverage LLMs’ pattern completion strengths while addressing their execution limitations through principled architectural modifications.
3 Structural Limits on Symbolic Computation
The computational split-brain syndrome manifests most clearly in symbolic computation, where LLMs can perfectly explain algorithms they cannot reliably execute. This systematic dissociation—drawing analogy to neurological conditions where different brain systems cannot coordinate effectively—emerges as geometric separation in representational space: models develop distinct pathways for “knowing about" procedures versus “executing” them.
This section examines the architectural constraints that prevent reliable symbolic reasoning through a causal chain of three interdependent factors. Importantly, no single constraint alone would be fatal—contextual averaging could be overcome with better training techniques, architectural limitations might be addressed through scaling, and instruction-execution disconnect could potentially be bridged through clever prompting. However, these three constraints reinforce each other systematically: contextual averaging prevents automatic domain binding (Claim 1), while architectural limitations make direct symbolic computation impossible through weight configuration alone, forcing models toward pattern storage (Claim 2), and next-token prediction treats algorithmic descriptions and execution traces as equivalent pattern-completion tasks, preventing instructional guidance from bridging this gap (Claim 3). Together, these interdependent factors reveal why LLMs function as sophisticated pattern completion engines rather than symbolic reasoners, with implications extending to relational reasoning as we explore in Section 4.
3.1 Claim 1: Contextual Averaging Prevents Automatic Domain Binding
Token embeddings in LLMs encode context-weighted averages that resist automatic domain binding, making stable symbolic circuits extremely difficult to form. Unlike traditional programs that explicitly bind variables to types and domains, LLMs derive token meanings from statistical patterns across all training contexts.
The Failure of Mathematical Binding.
The root of this problem lies in how LLMs learn token representations. During training, the model encounters “9.11” in countless different contexts—historical articles about September 11th, software version discussions, financial data, and mathematical expressions. The training objective forces the model to find a single embedding that works reasonably well across all these contexts.
Formally, this creates a fundamental tension. The training objective optimizes embeddings through next-token prediction:
$\mathcal{L}(\theta)=-\sum_{t}\log p_{\theta}(w_{t}\mid w_{<t})$ (1)
$e^{*}(w)=\arg\min_{e}\sum_{c\in C(w)}\mathbb{E}_{w^{\prime}\sim p_{\text{data}}(\cdot|c,w)}\left[-\log p_{\theta}(w^{\prime}\mid e,\;E(c\setminus w))\right]$ (2)
where $C(w)$ denotes all contexts containing token $w$. This produces context-averaged embeddings that trade off among multiple usage modes, similar to distributional semantics models like Word2Vec Mikolov et al. (2013).
In formal reasoning, we bind symbols to precise domains ($a=9.9,a\in\mathbb{R}$), stripping away non-mathematical associations. LLMs typically struggle to perform this domain binding reliably. When encountering “9.9”, they tend to preserve contextual associations rather than mapping to a clean mathematical representation.
To demonstrate this contextual contamination, we presented LLaMA2-7B-chat with prompts like “9.11 is a” and “9.9 is a”. The completions reveal dramatically different semantic associations:
  • For “9.11 is a”, the model produces: “day”, “date”, “remembrance”, “tragedy”
  • For “9.9 is a”, the model produces: “number”, “decimal”, “perfect”, “high”
This divergence shows “9.11” retaining historical event associations while “9.9” behaves as a scalar quantity. Quantitative analysis using negative log-likelihoods across different templates (DATE, VERSION, MEASURE) confirms these contextual biases. The missing operation is domain binding—establishing “9.11” as an element in $\mathbb{R}$ governed by mathematical laws rather than historical associations.
Failure of Embedding Geometry.
Symbolically stable circuits require representation spaces that preserve consistent structural relationships. While this doesn’t demand strictly Euclidean geometry, it requires systematic, predictable structure. As Equation 2 predicts, contextually-averaged embeddings resist the domain binding required for stable symbolic circuits.
Figure 1: Numeric embeddings in LLaMA2-7B-chat fail to preserve symbolic geometry: (a) irregular cosine distances between digit tokens; (b) bifurcation at “10.0” due to tokenization boundaries; (c) large distances between equivalent representations like “5” and “five”.
The embedding analysis reveals the extent of this geometric chaos (Figure 1 ). What should be the mathematically ordered space of numbers instead resembles a scrambled puzzle, with qualitatively similar failures observed across Qwen and Mistral models:
  • Panel (a): Cosine distances between digit tokens “1” to “10” show irregular, asymmetric patterns. Notable violations of numerical ordering include “10” being closer to “1” (distance 0.20) than to “9” (distance 0.55), and “6” being closer to “10” (distance 0.59) than to adjacent digits like “7” (distance 0.40).
  • Panel (b): Decimal embeddings exhibit clear clustering by first digit rather than numerical proximity. The sharp visual boundaries separate 9.x values (top-left dark block), 10.x values (middle diagonal band), and 11.x–12.0 values (bottom-right region), revealing tokenization-driven rather than mathematically-principled organization.
  • Panel (c): Symbolic equivalents show systematic misalignment in embedding space. Ideally, we should observe a single diagonal of low distances where digit-word pairs (“1”/“one”, “2”/“two”, etc.) align. Instead, the heatmap reveals fragmented diagonal patterns with significant irregularities. While some digit-word pairs show reasonable proximity (e.g., “5” is relatively close to “five”), the embedding space exhibits systematic failures: a group of higher-numbered words like “ten”, “eleven”, and “twelve” show poor alignment with any digit tokens, and paradoxically, “5” appears closer to “fifteen” than many other digits are to their correct word equivalents. This demonstrates that the embedding geometry fails to consistently preserve mathematical equivalence relationships.
These geometric failures directly impede the formation of stable arithmetic circuits that could operate reliably across different input formats or numerical ranges. McLeish et al. (McLeish et al., 2024) demonstrate that altering input geometry alone can dramatically improve simple numerical reasoning, which might appear to contradict our contextual averaging argument. However, their approach involves training models exclusively on arithmetic problems, which inadvertently avoids the fundamental contextual averaging problem faced by real-world LLMs. Once embeddings are contaminated through multi-domain training—where tokens like “9.11” accumulate conflicting associations across historical, versioning, and mathematical contexts—this geometric solution breaks down. The McLeish results thus support rather than refute our analysis: they show embedding geometry matters precisely because real-world training makes it unsolvable.
But suppose we could somehow solve the representation problem—imagine LLMs had perfect, mathematically-structured embeddings where “9.9” and “9.11” occupied precisely the right positions in embedding space. Would this enable reliable symbolic computation? Unfortunately, a deeper architectural constraint ensures the answer is no. Even with ideal symbolic representations, the computational mechanisms of transformer networks face fundamental limitations that force them toward pattern storage rather than algorithmic execution.
3.2 Claim 2: Feed-Forward Networks Resort to Pattern Storage
Given contextually-averaged symbolic representations (Claim 1), feed-forward networks face an additional fundamental constraint: the architectural impossibility of implementing exact symbolic operations through weight configuration alone. With direct computation blocked, FFNs resort to residual fitting by memorizing patterns rather than implementing symbolic procedures, despite their theoretical computational capacity.
The Computational Burden Falls on FFNs.
The transformer architecture creates a clear computational division. While attention mechanisms perform complex contextual computations—processing relationships between operands through learned weighted combinations and integrating information across the evolving sequence representation—they operate within the constraint that they can only combine existing representations through weighted averages. The generation of precise computational results must ultimately occur through the feed-forward networks, which apply position-wise transformations to produce new content.
Architectural Impossibility of Exact Computation.
Feed-forward networks in transformers implement an expand-compress architecture first proposed in Vaswani et al. (2017). For arithmetic queries, the required computational results aren’t directly available in the attention-processed representations and must be computed—this is where the FFN serves as the primary computational engine.
This architecture can be understood as an implementation of the Universal Approximation Theorem (UAT) Hornik (1991): expanding to a higher-dimensional space (e.g., 49,152 dimensions in GPT-3) provides the width needed for universal function approximation, before compressing back to the model’s working dimensions (12,288). The fundamental operation follows:
$\text{FFN}(x)=\text{ReLU}(xW_{1}+b_{1})W_{2}+b_{2}$ (3)
Despite this theoretically universal approximation capacity, FFNs face a fundamental constraint for symbolic operations. The ReLU-based architecture implements piecewise linear functions, and as we prove formally in Appendix A for the case of multiplication, piecewise linear functions cannot implement exact symbolic operations requiring higher-order interactions through weight configuration alone. Similar constraints apply to division, exponentiation, and logical operations that require non-linear variable interactions over unbounded domains—while FFNs can approximate such operations arbitrarily well within any bounded region, symbolic computation requires exact results across potentially unbounded inputs, creating a fundamental architectural constraint that the expand-compress design cannot overcome through scale alone.
This impossibility extends to other exact symbolic operations, creating a fundamental architectural constraint that the expand-compress design cannot overcome through scale alone.
Why Pattern Storage Wins.
Each layer follows the same computational pattern where attention processes the evolving representations and FFN generates the residual contribution. For layer $l$, the computation proceeds as:
$h_{\text{attn}}^{(l)}$ $=x^{(l-1)}+\text{Attention}_{l}(x^{(l-1)})$ (4)
$\Delta h^{(l)}$ $=\text{FFN}_{l}(h_{\text{attn}}^{(l)})$ (5)
$x^{(l)}$ $=x^{(l-1)}+\Delta h^{(l)}$ (6)
The key insight is that each layer’s FFN faces the identical architectural constraint—none can perform exact symbolic computation. Instead, they learn to coordinate their residual contributions toward the target output through hierarchical pattern decomposition:
  • Layer 1: Attention recognizes basic task structure (‘×’ indicates multiplication, operands ‘23’ and ‘17’), FFN contributes initial pattern fragment $\Delta h^{(1)}$ that begins moving toward ‘391’
  • Layer 2: Attention processes representations enriched by Layer 1’s computational work, FFN refines the approximation with $\Delta h^{(2)}$ based on learned pattern fragments
  • Layer L: Attention integrates all accumulated computational history, FFN makes final pattern-based adjustment $\Delta h^{(L)}$ toward the target
The final output becomes:
$\text{output}=\text{embeddings}+\sum_{l=1}^{L}\Delta h^{(l)}$ (7)
Critically, each FFN learns conditional pattern matching over attention-processed representations: “when attention gives me this enriched computational state $\to$ I output this residual contribution.” No individual layer performs algorithmic computation—each contributes learned pattern fragments that collectively approximate the target through residual accumulation.
Since exact algorithmic implementation remains architecturally impossible at each layer, gradient descent naturally converges toward hierarchical pattern assembly across the residual stream. This explains why models achieve high accuracy on familiar arithmetic patterns while failing systematically when problem structures deviate from training distributions—the system approximates multiplication through coordinated pattern retrieval rather than principled computation, leading to brittle performance on novel configurations with drops of up to 80% Srivastava et al. (2024); Mirzadeh et al. (2024).
Experimental Validation of Hierarchical Pattern Assembly.
To test this hierarchical pattern decomposition hypothesis, we analyzed layer-by-layer hidden state trajectories in LLaMA-2-7B during arithmetic computation [^1] Code and data for reproducing all embedding analyses and layer-by-layer convergence experiments will be made available upon publication.. For prompts like “23 × 96 =” and “63.7 + 3.5 =”, we measured how each layer’s representation at the prediction position converges toward the final layer’s output representation.
Figure 2: Experimental validation of hierarchical pattern assembly in arithmetic computation. Layer-by-layer convergence analysis for multiplication (left) and addition (right) shows progressive refinement from initial similarity of $\sim$0.07 to late-layer similarity of $\sim$0.48. Error bars represent standard error across 10 examples per operation. The consistent improvement patterns (0.404 for multiplication, 0.407 for addition) provide direct neural evidence for residual pattern fitting rather than algorithmic computation. Shaded regions indicate early (blue), middle (orange), and late (green) processing phases.
Figure 2 demonstrates strong empirical support for our theoretical framework. Both multiplication and addition exhibit nearly identical progressive refinement patterns: representations begin with low similarity to the final output ($\sim$0.07) and gradually converge through three distinct phases. This layer-by-layer analysis confirms that since each individual FFN faces the same architectural impossibility of exact computation, the only viable learning strategy becomes hierarchical pattern decomposition, where multiple FFNs coordinate their residual contributions through learned pattern sequences rather than implementing systematic algorithms.
This pattern storage behavior aligns with recent findings by Dziri et al. Dziri et al. (2023), who demonstrate that transformers memorize computation patterns from training rather than learning generalizable algorithms, with models succeeding primarily when test patterns appeared frequently in training data. This frequency-dependent behavior also explains some relational reasoning failures we examine in Section 4.
But this raises a puzzling question: if LLMs cannot execute symbolic operations reliably, why do they excel at explaining algorithmic procedures? When asked “How do you multiply two numbers?", models provide textbook-perfect explanations of the algorithm. This suggests a potential solution: perhaps these articulated procedures could implicitly guide execution, with the model’s comprehensive knowledge of algorithms compensating for its execution limitations.
Unfortunately, this intuitive hope fails due to a third architectural constraint that prevents instructional knowledge from rescuing symbolic computation.
3.3 Claim 3: Next-Token Prediction Decouples Instruction from Execution
The next-token objective treats algorithmic descriptions and execution examples as equivalent prediction tasks, preventing the binding of procedural knowledge to computational implementation. Even if architectural mechanisms existed to enable instructional guidance, the geometric separation of instruction and execution pathways ensures that procedural knowledge cannot rescue symbolic computation.
The False Promise of Instructional Binding.
Training data contains examples that might appear to support instruction-execution binding. Consider instructional sequences like:
“To multiply 23 × 17: First multiply 23 × 7 = 161, then multiply 23 × 10 = 230, finally add 161 + 230 = 391.”
Such examples seem to demonstrate algorithmic steps paired with execution traces, potentially enabling models to bind descriptions to implementations. However, next-token prediction processes these as sequential pattern completions without distinguishing instructional content from computational steps.
The expectation that exposure to such instruction-execution pairs would automatically enable binding represents a fundamental illusion about neural learning. Training models to recite algorithmic procedures does not automatically configure weights to perform those operations or enable inference-time binding between instruction and execution. The next-token objective does not provide a mechanism for such binding: it treats “how to multiply” and “what is 23×17” as independent text completion tasks.
Indiscriminate Processing Across Context Types.
The training corpus contains arithmetic in diverse contexts: instructional examples with step-by-step explanations, homework solutions, calculations embedded within articles, and standalone problem-answer pairs. Next-token prediction treats all of these—instructional descriptions, execution traces, and standalone calculations—as equivalent token sequences to predict, without distinguishing their functional roles.
This creates two separate learning pathways with distinct representational signatures: models learn to recite algorithms by pattern-matching instructional texts, while learning to execute operations by retrieving stored pattern fragments from training examples. Critically, these competencies become geometrically separated in the model’s latent space—instructional knowledge and execution capabilities occupy different representational regions.
Experimental Validation of Geometric Separation.
Figure 3: t-SNE projection of instruction, symbolic execution, and word-form execution embeddings from LLaMA2-7B-chat, across 50 problems each in five arithmetic operations. Instructions (circles), symbolic expressions (squares), and word-form results (triangles) occupy geometrically distinct regions, despite describing the same operation. This illustrates that LLaMA2-7B-chat stores instructional knowledge and executional know-how in separate regions of its latent space.
If our hypothesis is correct, we should be able to directly observe this separation in the model’s representational space. When a model encounters “To multiply 56 and 76, first break one number into parts…” versus “56 × 76 = 4256,” are these processed by the same neural pathways or completely different ones?
To test this, we constructed a dataset of 250 arithmetic problems spanning five operations, each expressed in three distinct forms:
  • an instructional sentence that explains the procedure (e.g., “To multiply 56 and 76, first break one number into parts…”),
  • a symbolic execution form (e.g., “56 $\times$ 76 = 4256”), and
  • a word-form result (e.g., “A worker makes 56 parts per day for 76 days, totaling 4256 parts”).
We embedded each sentence using LLaMA2-7B-chat and projected the resulting vectors with t-SNE. As shown in Figure 3, the model organizes these representations into clearly separated geometric clusters: instruction embeddings are well-separated from execution forms, and word-form results cluster more closely with symbolic executions than with instructions. This visualizes the structural dissociation between verbalized procedures and learned execution templates in the model’s latent space.
Figure 4: Cosine distances between cluster centroids for each operation-role pair, grouped by role: instructional texts (top block), execution strings (middle), and worded executions (bottom). Each centroid represents the average embedding of 50 samples per category, computed using LLaMA2-7B-chat. Clear geometric separation is observed between instructional and execution-related forms, with execution and worded execution clusters being closer to each other than either is to instruction. This provides quantitative support for the hypothesis that instruction-following and execution are learned as distinct representational pathways within the model’s embedding space.
We computed the mean embedding (centroid) for each (operation, role) cluster based on 50 examples per category. We then calculated pairwise cosine distances between these centroids and visualized the result as a heatmap (Figure 4). The cluster labels are grouped first by role (instruction, execution, worded execution) and then by operation. The plot reveals a clear block structure: instructional forms form one cluster with tight internal cohesion but are distant from both executional forms. In contrast, execution and worded execution are notably closer to one another, reflecting their shared emphasis on output form rather than procedural structure.
Inference-Time Dissociation.
At inference time, different prompts steer the model toward geometrically distinct subspaces in its representational space. When prompted for algorithmic explanation (e.g., “How do you multiply two numbers?"), the model gravitates toward the instructional subspace, retrieving pedagogical patterns from its training distribution. When prompted for calculation (e.g., “What is 56 × 76?"), the model seeks out the execution subspace, accessing stored arithmetic patterns.
These operate as independent systems developed through separate pattern-matching processes during training, explaining why models can fluently articulate multiplication procedures while failing to execute them reliably. The geometric separation we observed in Figure 4 manifests functionally as this prompt-dependent subspace selection, where the model’s response pathway is determined by which representational region the prompt activates rather than any principled binding between instruction and execution.
With all three architectural constraints now established—unstable symbolic representations, computational impossibility of direct symbolic operations, and geometric separation of instruction from execution—we can see how they interact to create the computational split-brain syndrome. No single constraint alone would be fatal, but together they systematically prevent LLMs from bridging the gap between comprehension and competence.
3.4 The Broader Pattern: From Architecture to Cognition
With all three escape routes blocked—unstable symbolic representations, architectural impossibility of direct computation, and geometric separation of instructional knowledge from execution—LLMs inevitably resort to sophisticated pattern storage and retrieval systems. This explains the computational split-brain syndrome: models become excellent tutors who can articulate multiplication procedures (through instructional pattern-matching from Claim 3) while executing them through brittle hierarchical pattern retrieval rather than principled computation.
Contrast with Symbolic Systems.
Traditional symbolic computation systems avoid these limitations through explicit design choices that transformer architectures lack:
  • Type Systems: Variables are bound to specific domains (e.g., $a\in\mathbb{R}$) with operations defined over those types, preventing contextual contamination.
  • Algorithmic Implementation: Operations like multiplication are implemented as explicit procedures rather than learned approximations.
  • Compositional Binding: Complex expressions maintain consistent variable bindings through syntactic structure.
Path-Dependent Fragmentation and Interpretability Challenges.
The residual fitting process is highly sensitive to training dynamics across diverse domains, suggesting why mechanistic interpretability findings often lack generalizability. Since arithmetic represents just one of many competing objectives, the order and context of examples during training significantly impact which pattern clusters emerge and where they are encoded. Training data ordering, scheduling decisions, and model architecture jointly determine which statistical regularities crystallize into seemingly coherent “circuits.”
Recent detailed mechanistic analysis confirms this training-dependency hypothesis. Nikankin et al. Nikankin et al. (2024) provide valuable evidence that arithmetic computation relies on distributed “bags of heuristics” spread across multiple layers, rather than implementing dedicated algorithmic circuits for basic operations ($+$, --, $\times$, $\div$). Their neuron-level analysis reveals that no coherent computational circuits exist for these fundamental operations. Importantly, they find that models within the same training lineage (Llama3-8B vs Llama3-70B) share more similar heuristic patterns than models from different training backgrounds (Pythia-6.9B, GPT-J), yet even within this shared lineage, the models differ in the degree and sophistication of these patterns. This training-dependency gradient—where shared training yields similar but not identical mechanisms—confirms that apparent “arithmetic mechanisms” reflect training-specific statistical regularities rather than universal computational principles.
Reversing Dennett: Comprehension Without Competence.
This pattern inverts what philosopher Daniel Dennett observed in natural systems, where competence typically precedes comprehension (Dennett, 2017). Simple organisms demonstrate complex behaviors before developing explicit understanding. LLMs exhibit the reverse: sophisticated explanatory capabilities coupled with unreliable execution—comprehension without competence.
This reversal reveals LLMs as fundamentally different from both biological intelligence and traditional computational systems. They excel at linguistic pattern completion while generally lacking the grounded symbolic manipulation capabilities that such fluent descriptions would suggest.
These representational, computational, and learning constraints reveal why LLMs function as sophisticated pattern completion engines rather than symbolic reasoners. The computational split-brain syndrome emerges as a fundamental property of transformer architectures optimized for pattern completion rather than symbolic manipulation.
4 From Arithmetic to Relational Reasoning
If our architectural analysis is correct, the computational split-brain syndrome should not be limited to arithmetic. The same three constraints—contextual averaging, architectural impossibility of exact computation, and instruction-execution disconnect—should create identical failure patterns wherever systematic symbolic manipulation is required.
Relational reasoning provides the perfect test case. Like arithmetic, it demands automatic binding from natural language to symbolic structure, followed by systematic execution of transformations. Just as “9.11” must be bound as a numerical quantity rather than a historical date, “Alice” must be bound to appropriate relational roles before applying logical transformations like $\forall x,y$ [Parent$(x,y)\rightarrow$ Child$(y,x)$].
Table 1: Two-Phase Computational Requirements. Both arithmetic and relational reasoning require automatic binding from natural language followed by systematic execution of transformations.
Phase Arithmetic Operations Relational Reasoning
Input Binding Automatically infer and bind: “9.11” $\rightarrow$ numerical value, “9.9” $\rightarrow$ numerical value Automatically infer and bind: “Alice” $\rightarrow$ PERSON, family relationships from syntax
Computational Execution Execute bound operations: $9.9>9.11$ Execute inference rules: Parent$(X,Y)\rightarrow$ Child$(Y,X)$
Consistency Maintain across multi-step procedures: $a>b\land b>c\rightarrow a>c$ Maintain variable bindings across multi-step inference: Parent$(A,B)\land$ Parent$(B,C)\rightarrow$ Grandparent$(A,C)$
Table 1 illustrates these parallel computational demands across domains. Both require the same fundamental operations: automatic symbolic binding from natural language, reliable rule execution, and consistency maintenance across multi-step reasoning chains. The same three architectural constraints predict systematic failures in relational reasoning: contextual averaging prevents clean symbolic binding, instruction-execution disconnect separates rule articulation from application, and pattern storage leads to memorized templates rather than systematic variable manipulation.
This theoretical extension predicts LLMs will exhibit identical split-brain syndrome across domains—fluent explanation of logical principles coupled with unreliable execution. The evidence strongly supports this prediction.
4.1 The Reversal Curse and Alice Problem
Consider two cases where LLMs fail at drawing basic conclusions that appear deceptively simple. The Reversal Curse (Berglund et al., 2023) tests models on their ability to perform bidirectional inference: models trained only on statements like “Tom Cruise’s mother is Mary Lee Pfeiffer” are tested on the reverse direction “Who is Mary Lee Pfeiffer’s son?”—a symmetric transformation humans perform unconsciously. The Alice problem (Nezhurina et al., 2024) presents a different challenge: models given “Alice has 4 sisters and 1 brother” must answer “How many sisters does Alice’s brother have?” This requires rebinding Alice from her own perspective to her brother’s perspective while maintaining compositional consistency across family relationships.
Empirical Failures
The magnitude of these failures is striking. Even after fine-tuning on fictional statements like “Uriah Hawthorne is the composer of Abyssal Melodies,” models cannot answer “Who composed Abyssal Melodies?” revealing the same automatic binding failures we identified in arithmetic domains. Experiments show that models fine-tuned on 1,000 fictional “A is B” statements achieved near-perfect accuracy on forward retrieval but only 7% accuracy on reverse queries(Berglund et al., 2023). This failure persists in real-world scenarios: when tested on celebrity facts, GPT-4 correctly answers forward questions like “Who is Tom Cruise’s mother?” 79% of the time, compared to only 33% for reverse questions like “Who is Mary Lee Pfeiffer’s son?”(Berglund et al., 2023). The model’s log-probability for correct answers in the reverse direction shows no improvement over random baseline, indicating systematic rather than accidental failure.
The Alice problem (Nezhurina et al., 2024) presents equally dramatic breakdowns. Given the seemingly simple prompt “Alice has 4 sisters and 1 brother” and asked “How many sisters does Alice’s brother have?”, state-of-the-art models including GPT-4, GPT-4o, and Claude 3 Opus show “strong collapse of reasoning” across most tested models(Nezhurina et al., 2024). Despite the straightforward answer (5: Alice plus her 4 sisters), comprehensive testing across multiple prompt variations and at least 30 trials per model revealed frequent failures and extreme performance fluctuations on trivial problem variations. When confronted with more complex family structures (AIW+), performance collapsed to “close to 0” even for the most capable models(Nezhurina et al., 2024). Notably, models exhibit strong overconfidence in wrong solutions while providing “confabulation-like” explanations that sound plausible but demonstrate fundamental failures in compositional reasoning and variable binding.
Computational Split-Brain in Relational Reasoning
These failures exemplify computational split-brain syndrome in relational reasoning, directly predicted by our architectural analysis in Section 3. Models demonstrably know both generic rules (“If X is Y’s mother and Y is male, then Y is X’s son”) and specific facts (“Tom Cruise’s mother is Mary Lee Pfeiffer”), yet systematically fail to execute the algorithmic steps required for inference. This reflects the same instruction-execution disconnect identified in Section 3.3, where procedural knowledge and computational implementation occupy geometrically separated representational spaces.
Consider the computational requirements for the reversal curse: given “Tom Cruise’s mother is Mary Lee Pfeiffer” and asked “Who is Mary Lee Pfeiffer’s son?”, models must automatically infer that the tokens represent PERSON entities, bind the implicit MOTHER_OF relationship from surface syntax, recognize the query seeks the inverse relationship, apply the symmetric transformation Mother$(x,y)\rightarrow$ Son$(y,x)$ in reverse, bind variables $x=$Mary Lee, $y=$Tom, and conclude Tom is Mary Lee’s son. Similarly, the Alice problem requires binding family structure from “has 4 sisters and 1 brother,” then rebinding Alice from SISTER role to SIBLING role to execute the perspective-shift computation.
Both tasks demand the symbolic manipulation capabilities that our architectural analysis shows current LLMs lack: contextual averaging prevents the clean symbolic binding required for role assignment, instruction-execution disconnect separates relational rule articulation from application, and pattern storage leads to memorized relationship templates rather than systematic variable manipulation.
Instead, models resort to directional pattern memorization. The reversal curse occurs because “A is B” patterns vastly outnumber “B is A” patterns in training corpora, causing asymmetric learning of fundamentally symmetric relationships. The same architectural constraints that prevent stable arithmetic circuits also undermine the variable binding and systematic transformation required for relational reasoning, creating the identical split-brain syndrome across domains—fluent explanation of logical principles coupled with unreliable execution.
We note that the reversal curse cannot be addressed through training data frequency balancing—simply ensuring equal exposure to “A is B” and “B is A” patterns. The issue is architectural: models learn separate representational pathways for each syntactic form rather than understanding the symmetric logical relationship. Even with perfect frequency balance, the system would still develop distinct pattern-matching rules instead of unified bidirectional binding. Solutions like LoCo-LM (Calanzone et al., 2025) succeed precisely because they bypass this limitation through explicit logical constraint enforcement in the training objective, essentially forcing systematic truth table coverage rather than relying on natural pattern completion tendencies.
4.2 Rule Applications, Logical Inconsistencies and Operator Failures
In contrast to the implicit complexity of the Reversal Curse and Alice problem, logical inconsistency tasks present LLMs with explicit formal structure that should simplify reasoning. These problems provide clear premises, stated rules, and well-defined logical relationships—yet LLMs continue to fail systematically. This failure despite explicit scaffolding reveals that the architectural constraints operate independently of how logical structure is presented.
Large language models exhibit a range of logical inconsistencies, as identified in LoCo-LM (Calanzone et al., 2025) and summarized by Cheng et al. (2025):
  • Negation Inconsistency: The model affirms both “X is an organism” and “X is not an organism.”
  • Implication Failure: Given “All birds are animals” and “An albatross is a bird,” the model fails to infer “An albatross is an animal.”
  • Reverse Implication Failure: From “If made of metal, then conducts electricity” and “X does not conduct electricity,” the model fails to infer “X is not made of metal.”
  • Deductive Chain Breakdown: Given “Nails are made of iron,” “Iron is a metal,” and “Metals conduct electricity,” the model fails to infer “Nails conduct electricity.”
These failures persist even when facts and rules are fully spelled out (e.g., “All birds can fly” and “Tweety is a bird” $\rightarrow$ “Tweety can fly”)—a setup that, unlike the Reversal Curse or Alice problem, explicitly states the binding relationships that should enable deduction if the architecture supported it.
(Vashishtha et al., 2025) demonstrates success on causal reasoning tasks that test rule applications without variable binding. Pattern matching works when models are trained from scratch on clean synthetic data, with a 67M-parameter transformer outperforming billion-parameter LLMs on causal axioms. This confirms that direct rule application falls within transformers’ capabilities when training conditions isolate logical operations from the contextual contamination of multi-domain pretraining.
Systematic Evaluation of Relational Reasoning
The comprehensive study by Li et al. (2024) provides decisive evidence of LLMs’ architectural limitations in relational reasoning. They systematically evaluated state-of-the-art LLMs on inductive logic programming (ILP) tasks across both natural language and truth-value matrix formats, comparing performance against Differentiable Logic Machines (DLM) (Zimmer et al., 2023)—specialized neural program induction models designed explicitly for logical reasoning.
The results are striking: despite being 100,000 times larger than DLM models, LLMs performed dramatically worse across all logical reasoning tasks. For family tree reasoning tasks requiring multi-step inference:
  • Simple relations (HasFather): LLMs achieved 47-100% accuracy vs. DLM’s perfect 100%
  • Complex relations (IsUncle): LLMs dropped to 0-49% accuracy vs. DLM’s 85%
  • Hierarchical relations (IsMGUncle): LLMs achieved only 0-48% vs. DLM’s 55%
Critically, these failures occurred across both natural language prompting and truth-value matrix representations. When given explicit logical structure through truth-value matrices—the same format used by specialized logical reasoning systems—LLMs still failed systematically. This definitively rules out training data scarcity or prompt engineering as explanations.
Hallucination in Logical Reasoning
The study revealed systematic hallucination patterns that illuminate the underlying architectural problems. LLMs generated contradictory reasoning processes, such as claiming “P8 is the father of P3 and also the mother of P3” when given explicit family relationships. These aren’t random errors but systematic pattern completion mistakes where models correctly identify surface facts but fail at the transformations required for compositional reasoning.
4.3 The Pattern Across Relational Reasoning
Our analysis reveals a convergent pattern across complexity levels: LLMs fail systematically regardless of how relational reasoning tasks are presented. Problems that appear simple (Reversal Curse, Alice problem) actually require sophisticated computational transformations that LLMs cannot execute. Problems that appear complex but provide explicit logical structure (truth-value matrices, clear logical relationships) should be easier to solve, yet LLMs still fail systematically. Table 2 summarizes the magnitude of these systematic failures, revealing consistently poor performance despite the apparent simplicity of many tasks.
Table 2: Systematic Failures in Relational Reasoning
Task Domain LLM Performance Comparison
Reversal Curse (forward) 79%
Reversal Curse (reverse) 7% Random baseline
Alice Problem (complex) $\sim$0%
Family Relations (IsUncle) 0-49% DLM: 85%
Family Relations (IsMGUncle) 0-48% DLM: 55%
This convergence is theoretically significant. If architectural limitations were merely about training data scarcity or prompt engineering, we would expect LLMs to succeed when given explicit logical structure. Instead, failure persists regardless of presentation format, confirming that the constraints operate at the architectural level—the same three limitations that prevent stable arithmetic circuits also undermine relational reasoning across all contexts.
Lifted Reasoning as the Missing Capability
The consistent failure across both implicit and explicit reasoning tasks points to a fundamental missing capability: lifted reasoning. What these tasks demand—and what LLMs fundamentally lack—is the ability to apply general rules over arbitrary entities rather than memorizing instance-specific patterns. Lifted reasoning underlies human-like generalization in logic, mathematics, and programming. Unlike pattern completion, which operates over specific instances, lifted reasoning requires variable abstraction (binding entities to abstract roles that can be systematically manipulated), rule generalization (applying logical transformations that work across arbitrary entity instantiations), and compositional consistency (maintaining variable bindings across multi-step inference chains).
Current transformer architectures lack the representational and computational mechanisms necessary for these operations. Models specifically designed for these capabilities, like Neural Logic Machines (NLM) (Dong et al., 2019) and Differentiable Logic Machines (DLM) (Zimmer et al., 2023), achieve lifted reasoning through explicit tensor representations for predicates and specialized circuits for logical operations. These architectures explicitly address the binding problem by representing relationships as tensors and implementing differentiable operations that maintain variable consistency—architectural innovations absent from standard transformers.
The performance gap persists even when LLMs are given identical inputs to specialized models: despite being 100,000 times larger than DLM models, LLMs performed dramatically worse across all logical reasoning tasks (Li et al., 2024), demonstrating that architectural constraints, not data availability, prevent reliable logical reasoning. This provides definitive evidence that the failures reflect fundamental limitations in how transformers process symbolic relationships, confirming our architectural analysis from Section 3.
This lifted reasoning gap explains why LLMs exhibit sophisticated pattern completion capabilities while systematically failing at the variable binding and rule application required for reliable relational reasoning. The computational split-brain syndrome manifests regardless of how logical structure is presented, confirming that the constraints operate at the architectural rather than training or prompting level. Understanding this fundamental limitation clarifies both what current LLMs can achieve as pattern completion engines and what alternative approaches might be needed to transcend these constraints.
5 Compensatory Strategies and Their Execution Gap
The computational split-brain syndrome reveals a fundamental paradox: LLMs can articulate principles they cannot reliably execute. However, this very comprehension capability suggests compensatory strategies. If models have memorized algorithmic procedures well, they can leverage these memorized patterns to unroll step-by-step solutions (self-scaffolding). If they can recognize problem types reliably, they can delegate to specialized external systems (tool delegation). Finally, rather than working around execution limitations, we can address them architecturally by integrating dedicated symbolic modules (hybrid architectures); essentially an internal tool-calling for critical operations.
Each approach leverages LLMs’ strengths while addressing execution limitations differently. Yet as we demonstrate, all three strategies converge on the same fundamental requirement: coordinating these approaches demands reliable metacognitive capabilities that current architectures lack (see Section 5.4).
5.1 Self-Scaffolding: Leveraging Comprehension for Step-by-Step Execution
Self-scaffolding exploits LLMs’ demonstrated ability to articulate memorized algorithmic procedures by having them generate explicit step-by-step decompositions, then attempt to execute each step themselves. This approach converts implicit reasoning into structured text patterns that leverage models’ pattern completion strengths.
Examples Across Domains.
A self-scaffolding approach to the problems of multi-step arithmetic, reversal curse, and Alice problem would generate:
“To multiply $742\times 89$: First, multiply $742\times 9=6678$. Then multiply $742\times 80=59360$. Finally, add $6678+59360=66038$.”
“Since Tom Cruise’s mother is Mary Lee Pfeiffer, and the relationship ‘mother of’ is symmetric to ‘son of’, Mary Lee Pfeiffer’s son must be Tom Cruise.”
“Alice has 4 sisters and 1 brother. From her brother’s perspective, he has Alice plus her 4 sisters, making 5 sisters total.”
The Illusion of Thinking study (Shojaee et al., 2025) is a recent example, demonstrating this pattern across planning puzzles, where Large Reasoning Models (LRMs) including Claude 3.7 Sonnet, DeepSeek-R1, and OpenAI’s o3-mini generate detailed algorithmic traces for problems like Tower of Hanoi while attempting execution.
5.2 Tool Delegation: Bypassing Model Execution Entirely
Tool delegation represents a fundamentally different strategy: rather than improving the model’s execution capabilities, it bypasses them entirely by delegating computational tasks to specialized external systems. This approach has been systematized in frameworks like ReAct (Yao et al., 2022), Toolformer (Schick et al., 2023), and various agent architectures that enable LLMs to interact with external APIs and computational tools.
Direct Delegation Examples.
For arithmetic, instead of attempting step-by-step execution, the model simply recognizes the problem type and delegates:
  • “This is a multiplication problem” $\rightarrow$ calculator.multiply(742, 89).
  • For database queries: “Find maximum sales” $\rightarrow$ SQL: SELECT MAX(sales) FROM data.
  • For logical reasoning: “Check satisfiability” $\rightarrow$ SMT_solver.check(constraints).
The comment on the Illusion of Thinking paper (Opus and Lawsen, 2025) provides a particularly striking example of tool delegation’s effectiveness. When Tower of Hanoi problems were reformulated as: “Solve Tower of Hanoi with 15 disks. Output a Lua function that prints the solution when called,” models achieved high accuracy across tested systems (Claude-3.7-Sonnet, Claude Opus 4, OpenAI o3, Google Gemini 2.5), completing in under 5,000 tokens. This dramatic improvement occurred because models generated algorithmic code rather than attempting to enumerate moves themselves. Similarly, Cheng et al. (Cheng et al., 2024) demonstrate that separating symbolic code generation from symbolic code execution improves deductive reasoning.
5.3 Hybrid Architectures: Specialized Modules for Symbolic Operations
Hybrid architectures represent a more fundamental response to the computational split-brain syndrome: addressing the limitations architecturally rather than through workarounds. These approaches integrate specialized computational modules within LLM systems to handle symbolic operations that current transformer architectures cannot reliably implement. Essentially, the idea is to mend the comprehension-competence gap internal to the model.
Domain-Specific Integration.
Several promising approaches have emerged across different symbolic domains. For arithmetic computation, OccamLLM (Dugan et al., 2024) and Internal General Computations (IGC) (Dietz and Klakow, 2024) incorporate specialized circuits that directly implement arithmetic operations, achieving perfect accuracy while avoiding pattern-based approximations. For logical reasoning, Logic-LM (Pan et al., 2023) integrates LLMs with symbolic solvers through translation-execution pipelines that maintain logical consistency.
Particularly encouraging are approaches like Differentiable Logic Machines (DLM) (Zimmer et al., 2023), which represent predicates as tensors and implement logic operations as differentiable functions. Unlike approaches that use neuro-symbolic losses to improve consistency in existing architectures, DLMs provide fundamental architectural support for rule learning and symbolic generalization, directly addressing the induction barriers we identified.
Functional Specialization Architectures.
Such integration, if done well, would combine pattern-matching flexibility with principled symbolic operations. Rather than emulating symbolic computation through pattern completion, such systems would implement dedicated mechanisms for variable binding, rule application, and systematic inference.
Current Mixture-of-Experts (MoE) architectures optimize primarily for computational efficiency—routing tokens to different parameter sets to increase model capacity—rather than functional specialization based on cognitive capabilities. A neuroscience-inspired approach could allocate specific experts to handle symbolic operations, relational reasoning, pattern completion, and other specialized functions. Just as the human brain has dedicated regions for executive reasoning (prefrontal cortex) and memory (temporal lobe), functionally partitioned MoE could develop experts that excel at exactly the symbolic binding and manipulation operations that current general-purpose feed-forward networks struggle to implement.
5.4 Coordination and The Metacognitive Bottleneck
Complementary Strengths and Shared Limitations.
The three compensatory strategies form a complementary toolkit for addressing the computational split-brain syndrome, each leveraging LLMs’ pattern completion strengths while circumventing execution limitations in different ways. Self-scaffolding requires no external components or architectural modifications, working entirely within existing capabilities by converting implicit reasoning into explicit textual decompositions. It excels in medium-complexity scenarios where individual steps fall within reliable pattern storage, as demonstrated by Large Reasoning Models successfully solving problems like the Alice puzzle where base LLMs fail (Mitchell, 2025). Tool delegation converts architectural impossibilities into architectural strengths, achieving perfect reliability by offloading computation to dedicated symbolic processors. When appropriate tools exist, this approach works reliably across complexity levels for well-defined domains while minimizing computational overhead.
However, each strategy faces critical limitations that reveal a deeper architectural constraint. Self-scaffolding incurs massive computational overhead (multiplying two 10-digit numbers requires approximately $10^{2}$ FLOPs on a CPU but $10^{12}$ FLOPs in transformer inference), suffers from planning failures when decomposition relies on unreliable pattern completion, and encounters execution unreliability when individual steps require symbolic manipulation beyond pattern storage capabilities (Shojaee et al., 2025). Tool delegation faces problem recognition requirements (identifying when and which tools to use), tool availability constraints (limited to well-defined domains with existing specialized systems), and composition challenges when complex problems require coordination between multiple tools or combinations of delegation and direct reasoning.
While hybrid architectures offer the most principled solution by mending the comprehension-competence gap internally through dedicated symbolic processing capabilities, they face fundamentally similar limitations to external tool delegation. Just as external tool calling requires recognizing when and which tools to use, hybrid architectures must coordinate between symbolic and pattern completion modules—essentially internal tool routing decisions. This introduces integration complexity, neural adaptation requirements that may constrain symbolic expressiveness, and the same metacognitive challenge of determining when symbolic processing versus pattern completion is appropriate. Moreover, the internal nature of this coordination makes it potentially harder to monitor, debug, or override compared to external tool delegation.
The Metacognitive Dependency.
Most critically, all three approaches converge on the same fundamental requirement: they demand reliable metacognitive assessment to coordinate when and how to apply each strategy. Self-scaffolding requires knowing when decomposition will help versus hurt, tool delegation requires recognizing appropriate problem types and available tools, and hybrid architectures must coordinate between symbolic and pattern completion modules. This shared dependency exposes a recursive problem: the compensatory strategies themselves require the very introspective abilities that the computational split-brain syndrome prevents.
Algorithm 1 Idealized Metacognitive Control Loop
1:  $P\leftarrow$ input problem
2:  $d\leftarrow$ assess_difficulty$(P)$requires introspection}
3:  if $d\leq\tau$ then
4:     return pattern_complete$(P)$
5:  else
6:     $P^{\prime}\leftarrow$ decompose_or_delegate$(P)$tool call / planner}
7:     return solve$(P^{\prime})$
8:  end if
However, compensatory strategies require careful coordination and expose deeper introspection requirements. Consider an idealized metacognitive control loop (Algorithm 1) as an illustrative framework: while this process could theoretically be executed through pattern completion—LLMs can perform individual steps like tool calling and decomposition using their existing capabilities (e.g., GPT-4’s tool-calling demonstrations)—it reveals fundamental challenges in metacognitive assessment. The difficulty assessment in line 2 and decomposition choice in line 6 would require metacognitive capabilities that face inherent limitations:
  • Complexity assessment (line 2): “Is this problem computationally hard for my specific architecture?” (But how can the model assess difficulty without attempting the task?)
  • Capability introspection (line 2): “Can I execute this operation reliably through my pattern storage?”
  • Tool appropriateness (line 6): “Do I have the right external system for this problem type?”
  • Decomposition quality (line 6): “Will breaking this into sub-steps actually help?” (The decomposition process itself can be brittle, as suggested by reasoning model studies in (Shojaee et al., 2025))
  • Step reliability (line 6): “Can I trust each sub-computation in my scaffolding?”
The deeper issue is metacognitive: while commercial LLMs have developed sophisticated capabilities for recognizing when to delegate computational tasks to external tools, they lack reliable self-awareness of their internal computational limitations when such delegation is unavailable. The same instruction-execution disconnect that prevents reliable symbolic computation also creates fundamental challenges for model self-assessment of their internal pattern completion capabilities. Recent work has identified these metacognitive limitations across LLMs, with systematic overconfidence and poor calibration, particularly on out-of-distribution tasks (Geng et al., 2024; Kadavath et al., 2022). This metacognitive deficit manifests critically in high-stakes applications: studies of medical reasoning show that LLMs “lack essential metacognition for reliable medical reasoning,” consistently failing to recognize knowledge limitations (Griot et al., 2025).
To test the importance of metacognition, we designed a tightly controlled experiment using $n$-digit × $n$-digit multiplication tasks with place-value decomposition and simulated column-wise addition, and tested Claude Sonnet 4, GPT-4o, and Gemini 2.5 Flash across three conditions:
  • Zero-shot direct calculation: Models were asked to calculate multiplication problems directly with explicit constraints against scaffolding, decomposition, or external tools (“Calculate this multiplication using only your internal reasoning, without using any tools or code”). All models achieved 0% accuracy on 10-digit problems, demonstrating universal computational split-brain syndrome when constrained to pure transformer computation. At 5-digit complexity, Claude achieved 10.5% accuracy while GPT-4o and Gemini remained at 0%.
  • Self-generated decomposition: We tested whether models could independently generate appropriate decomposition prompts for multiplication problems. When asked to create step-by-step decomposition instructions, all models succeeded on 5-digit tasks. For 10-digit problems, GPT-4o produced prompts with 65% accuracy in decomposition steps and 94.2% overall quality, while Claude Sonnet 4 and Gemini 2.5 Flash maintained perfect decomposition accuracy.
  • Golden decomposition execution: We tested execution using golden decomposition prompts that explicitly guide step-by-step computation using place-value breakdown. Models were constrained to use only internal capabilities without external tools or self-correction. Table 3 summarizes performance across two difficulty levels.
Table 3: Performance on multiplication problems with golden decomposition prompts
Complexity Metric GPT-4o Gemini 2.5 Claude Sonnet 4
5-digit Overall Accuracy 5% (1/20) 95% (19/20) 100% (20/20)
Step-wise 95–100% 95–100% 100%
Sum Errors 19/20 1/20 0/20
10-digit Overall Accuracy 0% (0/21) 0% (0/20) 0% (0/20)
Step-wise 76–100% 95–100% 95–100%
Sum Errors 18/21 20/20 20/20
With 5-digit numbers, GPT-4o achieved 95–100% accuracy on individual multiplication steps yet failed all but one problem due to arithmetic errors in the final summation. Gemini 2.5 Flash performed near-perfectly (95% overall accuracy), while Claude Sonnet 4 achieved perfect performance. As complexity increased to 10-digit numbers, all models degraded: GPT-4o’s step-wise accuracy collapsed to 76–100% with systematic errors, while Claude and Gemini maintained high step accuracy (95–100%) but failed all final answers due to summation errors.
Critically, when given natural prompts without explicit constraints, these models automatically delegate multiplication problems to external computational tools, achieving perfect accuracy through sophisticated tool-routing capabilities. Our experiments reveal what happens when this learned compensatory strategy is disabled, exposing the underlying computational split-brain syndrome that automatic tool delegation normally masks.
These results reveal a fundamental limitation: even with perfect algorithmic decomposition, neither individual computational steps nor aggregation operations remain reliable as per-step complexity increases. Models face a computational trilemma: either (1) execute multi-step procedures internally and risk systemic breakdowns, (2) rely entirely on external tools, or (3) decompose steps further—which introduces new failure points even under perfect prompting.
Real-world applications involve countless multi-step algorithms of unknown complexity where external computational verification is unavailable. This necessitates reliable metacognitive assessment of internal capabilities: models must recognize their computational limitations and coordinate appropriate compensatory strategies without external guidance. While commercial models have learned sophisticated tool delegation for well-defined computational tasks, they lack the self-awareness needed to assess the reliability of their internal pattern completion processes.
This raises a fundamental question about the nature of metacognition itself. Large Reasoning Models and approaches like Reflexion (Shinn et al., 2023) appear to learn metacognitive skills: knowing when to reflect, when to distrust their initial judgment, when to backtrack. During training, external verifiers provide ground truth signals about when reflection is needed, allowing models to learn patterns about failure recognition. At inference time, this “metacognition” manifests as pattern association: recognizing cues that previously correlated with the need for reflection. But is this genuine self-awareness or sophisticated pattern matching? The fundamental challenge remains that any metacognitive assessment—“I should doubt this answer,” “This reasoning seems flaky”—must itself emerge from pattern recognition over the same representational pathways we have shown to be unreliable for systematic reasoning.
Perhaps this mirrors human cognition itself—our sense of self-awareness may similarly emerge from pattern recognition over internal states rather than direct introspective access to our computational processes. This suggests that the distinction between “genuine” and “pattern-matched” metacognition may be less meaningful than initially apparent: if both humans and LLMs rely on learned associations to assess their own reasoning, the key question becomes not the authenticity of self-awareness, but its reliability and scope.
Summary
Understanding these metacognitive requirements clarifies a fundamental limitation of current approaches to the computational split-brain syndrome. While the three compensatory strategies form a powerful toolkit—each leveraging LLMs’ pattern completion strengths in different ways—they all converge on the same recursive problem: coordinating these sophisticated workarounds requires the very introspective abilities that the split-brain syndrome prevents.
Even with perfect algorithmic decomposition, as our experiments demonstrate, execution gaps persist at the individual step level. This suggests that the instruction-execution disconnect operates at a more fundamental architectural level than compensatory strategies can address. The metacognitive bottleneck reveals why current LLM capabilities, however sophisticated, remain fundamentally constrained: they excel as pattern completion engines but lack the self-awareness needed to reliably coordinate between comprehension and competence. This clarifies both what current architectures can achieve and what limitations persist even with the most sophisticated compensatory mechanisms.
6 Pattern Completion: Strengths and Limitations
Having examined LLMs’ systematic limitations in symbolic computation (Section 3), relational reasoning (Section 4), and compensatory strategies (Section 5), two critical questions emerge: How do we explain the remarkable power of LLMs as they are today? And what are the consequences of their architectural limitations? Our analysis points to pattern completion as both the source of their strengths and the root of their symbolic weaknesses.
Definition 1.
General intelligence = sophisticated pattern completion across diverse domains, with robust knowledge retrieval and flexible task adaptation. Generalizable intelligence = systematic rule discovery and principled reasoning that can be deployed to novel tasks and drives scientific progress.
We demonstrate that LLMs excel at the former but encounter fundamental barriers at the latter. The following analysis traces where today’s LLMs succeed on general intelligence but fail on generalizable intelligence.
Multi-level Pattern Completion.
LLMs achieve impressive capabilities through a hierarchy of pattern completion mechanisms. At the most basic level, next-token prediction enables fluent text generation, while higher up, the same mechanism supports chain-of-thought reasoning as pattern-matching operations over common reasoning structures.
Most recently, test-time computation allows these models to upgrade from LLMs to LRMs, functioning as self-programming metacomputers that dynamically adjust reasoning strategies based on task complexity. Problem decomposition, exploration, trial-and-verification, and backtracking can all be viewed as higher-level patterns for task execution.
This pattern-matching approach helps to explain why LLMs excel at tasks that can be solved through pure memorization (like retrieving facts) or tasks that appear to require symbolic manipulation but can actually be solved through pattern matching (like solving standardized math problems in familiar formats).
The Performance Cliff: From Pattern Recognition to Rule Discovery.
Because LLMs conflate similarity search with logical inference, recent empirical evidence from the ARC-AGI benchmark progression crystallizes the distinction between general and generalizable intelligence. ARC-AGI-1 evaluates pattern recognition and single-rule application in visual reasoning tasks, requiring models to identify and apply transformations from minimal examples (Chollet, 2019). A forthcoming benchmark extension (ARC-AGI-2) advances this evaluation by specifically targeting rule discovery and systematic application: tasks require inferring multiple interacting rules from limited observations, applying them compositionally across sequential steps, and adapting rule application based on contextual cues [^2] ARC-AGI-2 details from pre-release technical reports available at https://github.com/fchollet/ARC-AGI.
The results reveal a fundamental performance cliff: while OpenAI’s o3 achieved 87.5% on ARC-AGI-1 through sophisticated pattern completion, the same architecture scored $<3\%$ on ARC-AGI-2 (Chollet et al., 2025; ARC Prize Foundation, 2025). Performance plummets once pattern completion loses its near-neighbor anchors from training, despite identical instruction formats. This cliff occurs precisely where tasks transition from pattern recognition to genuine rule discovery and systematic reasoning—exposing the architectural limitations we have identified throughout this paper.
This performance cliff exemplifies our architectural analysis: LRMs achieve improvements through sophisticated self-scaffolding and test-time computation (compensatory strategies), but the underlying computational split-brain syndrome persists—they remain pattern completion engines that fail when tasks require genuine rule discovery and systematic reasoning beyond their memorized patterns.
Inductive Reasoning and the Limits of Generalizable Intelligence.
Inductive reasoning spans a spectrum of complexity, from simple function learning to discovering fundamental principles that govern entire domains. Consider the challenge of deriving an algorithmic solution from scratch. Recent work by Cheng et al. (2024) demonstrates that LLMs can excel at the simplest level—learning mathematical functions like base conversion from input-output examples—achieving near-perfect accuracy. At a higher level, FunSearch (Romera-Paredes et al., 2023) shows that LLMs can discover new algorithmic constructions for combinatorial problems through evolutionary code search, finding novel solutions to established mathematical challenges like the cap set problem.
However, discovery of novel algorithms such as solving the Tower of Hanoi problem from scratch requires a much stronger set of capabilities. The optimal algorithm that works for any number of disks $n$ requires recognizing that moving $n$ disks involves first moving $n-1$ disks, then the largest disk, then the $n-1$ disks again. This recursive insight, itself a meta-pattern, is even more challenging than what ARC-AGI-2 tests. Recent evaluation shows that state-of-the-art reasoning models like Claude 3.7 Sonnet and OpenAI’s o3-mini exhibit complete performance collapse beyond 8-9 disks, despite having sufficient token budget to continue (Shojaee et al., 2025).
Scientific discovery derives compact governing principles from noisy, real-world observations. Recent breakthroughs demonstrate that AI can indeed discover such principles, but through architecturally specialized approaches. AI Feynman succeeds through a physics-constrained symbolic search, whereas the LLM symbolic regression relies on pattern-based guessing (Udrescu and Tegmark, 2020; Makke and Chawla, 2022). These successes require domain-specialized architectures with strong inductive biases, such as geometric priors, physics constraints, and symbolic manipulation engines. It remains unclear whether pattern completion alone, however powerful, can accomplish genuine scientific discovery.
7 Conclusion
The apparent intelligence of LLMs belies a deeper fragility. Despite their success in pattern-rich tasks, these models consistently fail to generalize principles, execute symbolic computations, or reason reliably—even under idealized conditions. Our analysis reveals that these failures are not incidental but structural: LLMs dissociate instruction interpretation from execution, leading to a computational split-brain syndrome that prevents the formation of robust, composable operations.
This diagnosis clarifies why interpretability tools often uncover expressive artifacts rather than reliable algorithms, and why model self-explanations may diverge from the actual pathways used in computation. It also explains how training order can influence which internal patterns are embedded and where—even when outward behavior remains stable—causing interpretability efforts to surface path-dependent artifacts rather than consistent mechanisms. These insights reframe current conversations about emergence and scaling: more data or parameters cannot resolve bottlenecks rooted in architectural design.
If we seek genuinely general and generalizable intelligence, we must look beyond pattern completion engines. Future systems will need metacognitive scaffolding, lifted representations, and architectural support for principled execution. Our findings offer not only a critique of current LLMs, but a foundation for building the next generation of intelligent systems: one that can reason, not just react.
These findings have immediate implications for high-stakes applications where reliable reasoning is critical. Current LLMs require careful scaffolding, external verification, or hybrid architectures rather than deployment as standalone reasoning systems in domains like medical diagnosis, legal analysis, or safety-critical decision making. The computational split-brain syndrome we identify suggests that apparent fluency in explaining procedures should not be mistaken for reliable execution capability.
Appendix A Theoretical Proofs of FFN Limitations
A.1 Piecewise Linearity of Feed-Forward Networks
Theorem 1 (FFN Piecewise Linearity).
For any choice of weights and biases, the function implemented by a feed-forward network with ReLU activation is piecewise linear.
Proof.
Consider the FFN computation: $\text{FFN}(x)=\max(0,xW_{1}+b_{1})W_{2}+b_{2}$
  1. 1.
    Each neuron in the hidden layer computes $h_{i}=\max(0,x\cdot w_{i}+b_{i})$ where $w_{i}$ is the $i$-th column of $W_{1}$. The linear function $x\cdot w_{i}+b_{i}$ defines a hyperplane in the input space.
  2. 2.
    Each neuron’s ReLU activation partitions the input space based on $x\cdot w_{i}+b_{i}=0$, creating regions where the neuron is active ($x\cdot w_{i}+b_{i}>0$) or inactive ($x\cdot w_{i}+b_{i}\leq 0$).
  3. 3.
    The hyperplanes from all neurons intersect to partition the input space into polyhedral regions, each with a fixed pattern of active/inactive neurons.
  4. 4.
    Within any region, the activation pattern is constant, so the output is $\sum_{i\in\text{active}}(x\cdot w_{i}+b_{i})W_{2}[:,i]+b_{2}$. Since this is a linear combination of the input $x$, the function is linear within each region.
  5. 5.
    Therefore, the overall function is piecewise linear: linear within each region, with boundaries defined by the hyperplanes.
A.2 Impossibility of Exact Multiplication
Theorem 2 (Multiplication Impossibility).
No piecewise linear function can implement exact multiplication $f(x,y)=xy$ over an unbounded domain.
Proof.
Suppose, for contradiction, that some piecewise linear function $g$ implements exact multiplication over $\mathbb{R}^{2}$.
  1. 1.
    The multiplication function $f(x,y)=xy$ has mixed partial derivative $\frac{\partial^{2}f}{\partial x\partial y}=1$ everywhere.
  2. 2.
    Any piecewise linear function has zero mixed partial derivatives within each linear region (since linear functions have constant first derivatives and zero second derivatives).
  3. 3.
    For $g$ to equal $f$ everywhere, we would need $\frac{\partial^{2}g}{\partial x\partial y}=1$ everywhere, but this contradicts the piecewise linear property.
  4. 4.
    More constructively: consider any bounded region $R$ where $g$ is linear. Within $R$, we have $g(x,y)=ax+by+c$ for some constants $a,b,c$. But $ax+by+c\neq xy$ for all $(x,y)\in R$ unless $R$ contains at most finitely many points.
  5. 5.
    Since multiplication requires unbounded curvature (the function $xy$ becomes arbitrarily steep as $|x|$ or $|y|$ increases), no finite partition into linear pieces can approximate it exactly over an unbounded domain.
This impossibility extends to other non-linear operations requiring precise symbolic manipulation, establishing fundamental computational limits for FFN architectures in symbolic reasoning tasks.
Appendix B Metacognitive Assessment Experiments
B.1 Experimental Design
We tested three state-of-the-art models—Claude Sonnet 4, GPT-4o, and Gemini 2.5 Flash—on multiplication problems using “golden decomposition” prompts that provide explicit step-by-step procedural guidance. Models were constrained to use only internal capabilities without external tools or self-correction.
We evaluated two complexity levels: 5-digit numbers (range 10,000–99,999) and 10-digit numbers (range 1,000,000,000–9,999,999,999), with 20 problems per complexity level per model.
B.2 Sample Prompt
The following shows a representative prompt structure used in our experiments:
Problem: Calculate 93810 × 24592
Method: Use place value decomposition to break this into simpler steps.
Step 1 - Decompose 24592 into place values:
24592 = 2 + 90 + 500 + 4000 + 20000
Step 2 - Calculate each partial product:
Step 1: 93810 × 2 = ?
Step 2: 93810 × 90 = ?
Step 3: 93810 × 500 = ?
Step 4: 93810 × 4000 = ?
Step 5: 93810 × 20000 = ?
Step 3 - Add all partial products:
Sum: ? + ? + ? + ? + ? = ?
Final Answer: ?
This methodology eliminates planning uncertainty by providing the exact algorithmic steps, allowing us to isolate execution limitations from decomposition failures.
B.3 Implementation Challenges
GPT-4o exhibited systematic processing limitations: when given the full batch of 20 samples, it repeatedly failed to complete the task. When reduced to batches of 5 samples, GPT-4o often omitted intermediate calculation steps, proceeding directly to final answers without showing the required step-by-step work. This behavior required individual problem presentation to obtain reliable step-by-step responses.
Claude Sonnet 4 and Gemini 2.5 Flash processed full batches reliably and consistently followed the decomposition structure as specified.
B.4 Key Findings
The experiments revealed universal computational split-brain syndrome across all three model families. At 5-digit complexity, performance varied significantly: GPT-4o achieved 5% overall accuracy despite 95–100% step-wise accuracy, Gemini achieved 95% overall accuracy, and Claude achieved perfect performance.
At 10-digit complexity, all models exhibited complete breakdown (0% overall accuracy) but through different failure modes. GPT-4o showed degraded step-wise accuracy (76–100%), while both Gemini and Claude maintained excellent step accuracy (95–100%) but failed completely at final summation.
These results provide direct empirical evidence for the instruction-execution disconnect: models can follow algorithmic procedures perfectly yet systematically fail at computational implementation, demonstrating fundamental limitations in bridging comprehension and competence even under optimal procedural guidance.